9. Define elasticity of demand. Suppose, the demand function is Qd = 180 – 2P and supply function is Qs = 5+0.5P. Calculate the price elasticity of demand and supply. Calculate also consumer’s surplus and producer’s surplus.

14. Refer to question no. 9. If the government arbitrarily set the price $80, calculate fictional gain or loss of the industry. Calculate also the deadweight loss, consumer’s surplus and producer’s surplus.

15. Refer to question no. 9. If the government arbitrarily set the price $60, calculate fictional gain or loss of the industry. Calculate also the deadweight loss, consumer’s surplus and producer’s surplus. Note: If
Answer #1

Q9. At equilibrium, Demand is equal to the supply

So Q=180-2*70=180-140=40

The elasticity of Demand is

The elasticity of Supply is

The inverse demand function is given by

The inverse supply function is given by

Consumer surplus is

Producer Surplus is

Q.14.  At P=$80

Q=180-2*80=20

Loss of industry is 40-20=20

Deadweight loss is given by

Consumer Surplus is given by

Producer Surplus is given by

Q15.  Now P=$60

So Q=180-2*60=180-120=60

Gani in Industry 60-40=20

Deadweight Loss is given by

Consumer Surplus is given by

Producer Surplus is given by